2. You create a company selling printed shirts. You find that for a price of $ 11 for each shirt you manage to sell 1800 shirts a month. If the price of each shirt is $ 13, you can only sell 1450 shirts.

a. Assuming that the number of shirts sold can be modeled linearly based on the price of each shirt, determine the equation that models the total number of shirts sold in a month.

b. What does the slope mean?

c. You leave an employee in charge of selling the shirts. This puts a price and only managed to sell 1540 shirts. What was the price of the shirts?
Answer #1

1.
(Q-1450)=(1800-1450)/(11-13)*(P-13)
=>Q-1450=-175*(P-13)
=>Q=-175P+3725

2.
This means for 1 $ increase in price, quantity decreases by 175 units

3.
=(1540-3725)/(-175)
=12.48571429